# Hunting for extra dimensions in black hole shadows

### v1

@@ IV.1 Constraints on the $\mathrm{AdS_{5}}$ curvature radius
In our analysis, we will follow the methodology developed in (vagnozzi2022horizon, ), aiming to obtain bounds on the $\mathrm{AdS_{5}}$ curvature radius. Therefore, the observed angular radius – which has ring-like features in the horizon-scale EHT image of Sgr A* –, must be compared to the theoretically determined angular radius of the black hole shadow. From the precise determination of both mass and distance of Sgr A*, the central supermassive black hole of our galaxy, we can verify the accordance, within the uncertainty allowed, between the predicted theoretical results by this study and the empirical data found by the EHT observations (akiyama2022millimeter, ). On its turn, the mass and distance of Sgr A* have been studied over the last few decades by exploring the dynamics of star clusters and, in particular, the movement of stars close to the center of the galaxy. According to Ref. (vagnozzi2022horizon, ), the shadow radius measurements are restricted to the following regions determined by the average between the estimates based on the VLTI and Keck instruments, within $1\sigma$ and $2\sigma$ intervals, respectively,
$4.55\lesssim R_{s}/M\lesssim 5.22,\qquad\text{and}\qquad 4.21\lesssim R_{s}/M\lesssim 5.56$ (24)
@@
In Figure 4a, we show the expected deviations in the behavior of the shadow radius owing to the extra dimension. As we see, the increase in the $\mathrm{AdS_{5}}$ curvature radius $\ell$ results in a consequent increase in the radius of the black hole shadow. In our analysis, we have considered the average in the mass from KECK- and VLTI-based estimates. Therefore, from the empirical data obtained by observations of EHT, we can impose the limits on curvature radius at the $95\%$ confidence level (CL), such that $\ell\lesssim 6.5\times 10^{9}\,\mathrm{m}\simeq 4.3\times 10^{-2}\,\mathrm{AU}$, as shown in Figure 4b. At this point, we can compare our results with another which investigates similar physical systems. In this case, the absence of deviations on observation of the highly circular shadow of black hole M87* provides constraints on the $\mathrm{AdS_{5}}$ radius so that $\ell\lesssim 170\,\mathrm{AU}$ (Vagnozzi:2019apd, ). In turn, the first bounds obtained on the extra dimensions by the study of the time-lag between the detection of gravitational waves (GW) and electromagnetic (EM) signals from the merging of a binary system detected by the event GW170817 have established at $68\%$ CL an upper limit of $\ell\lesssim 0.535\,\mathrm{Mpc}\simeq 1.10\times 10^{11}\,\mathrm{AU}$ (Visinelli:2017bny, ), whereas our analysis found $\ell\lesssim 1.0\times 10^{-2}\,\mathrm{AU}$ for the same confidence interval. It is worth mentioning that, as we have seen, our study yields stronger constraints than those found by similar physical systems, according to the best of our knowledge. Ultimately, although we obtained weaker constraints compared to results obtained from sub-mm scale gravity experiments, this approach has set independent limits and can aid the investigation of the geometric structure of the universe.
@@
Figure 4: The left panel shows the changes in the behavior of the shadow radius with the variation of the $\mathrm{AdS_{5}}$ radius $\ell$. In turn, in the right panel, we found bounds from the direct comparison between the results obtained by our theoretical analysis with the EHT experimental data.
V Final Remarks
Recently, higher-dimensional theories have drawn increasing attention. As has been pointed out, the seek for evidence of large extra dimensions has been performed in several experimental researches, making its detection a challenging problem in itself. In this work, we have studied the quasinormal modes of a black hole in a brane-world model, aiming to search for traces of extra dimensions. By employing the WKB method, we compute the eikonal quasinormal frequencies. This complementary analysis also allowed us to confirm the theoretical prediction for the black hole shadow radius from the convergence between the WKB- and null geodesics-method estimates.
@@ V Final Remarks
From recent measurements of the shadow image of Sgr A* – a supermassive black hole in the center of our galaxy – we estimate deviations owing to an extra dimension in the RSII scenario on the theoretical shadow radius predicted. Given the inaccuracy involving the measured spin of Sgr A*, which may even be zero but still can provide compatible results with recent measures (Fragione:2020khu, ), we explore the possibility of describing them as a non-rotating RSII black hole. In this case, we have considered the Garriga-Tanaka solution that takes a classical approach to linearized brane-world gravity and describes a black hole solution in the weak field limit.
@@
As we saw, the curvature radius of the $\mathrm{AdS_{5}}$ spacetime is positively correlated to the shadow radius so that the increasing/decreasing in one leads to correspondingly increasing/decreasing in the other. Furthermore, we found that $\ell=0$, i.e., in the absence of new physics coming from extra dimension, one recovers the standard estimates expected for the Schwarzschild background. By investigating the null geodesic curves, we determine the shadow of the RSII black hole. From recent observational data of the shadow of Sgr A* released by the Event Horizon Telescope collaboration, one may impose constraints on the curvature radius of $\mathrm{AdS_{5}}$ spacetime. Therefore, we find more stringent limits on $\ell\lesssim 4.3\times 10^{-2}\,\mathrm{AU}$ (at $95\%$ CL) than previously obtained for similar physical systems, to the best of our knowledge. Moreover, non-zero predicted curvature radius also leads to agreement with the observed shadow radius within $68\%$ confidence interval, providing $\ell\lesssim 1.0\times 10^{-2}\,\mathrm{AU}$.
@@
We would like to highlight that although our analysis provides weaker bounds than those obtained from collider experiments or sub-mm precision gravity tests, which have set limits on the non-Newtonian force, in this study, we have provided conservative bounds since limits ourselves to investigate the shadow radius in a non-rotating RSII-type black hole background. Further, the bounds reported by us have been obtained in a manner complementary to the studies previously mentioned. Finally, the newly begun era of observational astronomy opens several avenues for future research that might be explored to obtain new constraints on the $\mathrm{AdS_{5}}$ curvature radius, for instance, or even to test extensions of general relativity, therefore shedding light into seeking for physics beyond the SM.
@@ Acknowledgements.
We would like to thank CNPq, CAPES and CNPq/PRONEX/FAPESQ-PB (Grant No. $165/2018$), for partial financial support. FAB acknowledges support from CNPq (Grant No. $309092/2022-1$). ASL acknowledges support from CAPES (Grant No. $88887.800922/2023-00$). JAVC would like to thank FAPESQ-PB/CNPq (Grant No. $077/2022$) for financial support.

### v2

@@ III Constraints on the $\mathrm{AdS_{5}}$ curvature radius
In our analysis, we aim to obtain bounds on the $\mathrm{AdS_{5}}$ curvature radius. Therefore, the observed angular radius—which has ringlike features in the horizon-scale EHT image of Sgr A*—must be compared to the theoretically determined angular radius of the black hole shadow. From the precise determination of both mass and distance of Sgr A*, the central supermassive black hole of our galaxy, we can verify the accordance, within the uncertainty allowed, between the predicted theoretical results by this study and the empirical data found by the EHT observations (akiyama2022millimeter, ). On its turn, the mass and distance of Sgr A* have been studied over the past few decades by exploring the dynamics of star clusters and, in particular, the movement of stars close to the center of the galaxy. According to Ref. (EventHorizonTelescope:2022xqj, ), the shadow radius measurements are restricted to the following region determined by the estimates based on the Keck and Very Large Telescope Interferometer (VLTI) instruments, within $1\sigma$ confidence interval, respectively,
$4.5\lesssim R_{s}/M\lesssim 5.5,\quad\text{and}\quad 4.3\lesssim R_{s}/M\lesssim 5.3$ (15)
In Fig. 3a, we show the expected deviations in the behavior of the shadow radius owing to the extra dimension. As we see, the increase in the $\mathrm{AdS_{5}}$ curvature radius $\ell$ results in a consequent increase in the radius of the black hole shadow. In our analysis, we have considered the mass from Keck- and VLTI-based estimates (vagnozzi2022horizon, ). Therefore, from the empirical data obtained by observations of EHT, we can impose the upper limits on curvature radius at the $68\%$ confidence level (CL), such that
$\ell_{\left(\text{Keck}\right)}\lesssim 5.7\times 10^{9}\,\mathrm{m}\simeq 3.8\times 10^{-2}\,\mathrm{AU},\quad\text{and}\quad\ell_{\left(\text{VLTI}\right)}\lesssim 2.4\times 10^{9}\,\mathrm{m}\simeq 1.6\times 10^{-2}\,\mathrm{AU},$ (16)
where $1\mathrm{AU}$ is one astronomical unit. These results are shown in Fig. 3b.
@@
At this point, we can compare our results with another which investigates similar physical systems. In this case, the absence of deviations on observation of the highly circular shadow of black hole M87* provides constraints on the $\mathrm{AdS_{5}}$ radius so that $\ell\lesssim 170\,\mathrm{AU}$ (Vagnozzi:2019apd, ). In turn, the first bounds obtained on the extra dimensions by the study of the time lag between the detection of gravitational waves (GW) and electromagnetic (EM) signals from the merging of a binary system detected by the event GW170817 have established at $68\%$ CL an upper limit of $\ell\lesssim 0.535\,\mathrm{Mpc}\simeq 1.10\times 10^{11}\,\mathrm{AU}$ (Visinelli:2017bny, ), whereas our analysis found $\ell_{\left(\text{Keck}\right)}\lesssim 3.8\times 10^{-2}\,\mathrm{AU}\left(\ell_{\left(\text{VLTI}\right)}\lesssim 1.6\times 10^{-2}\,\mathrm{AU}\right)$ for the same confidence interval. It is worth mentioning that, as we have seen, our study yields more stringent constraints than those found by similar physical systems, according to the best of our knowledge. Finally, although we obtained weaker constraints compared to results obtained from sub-mm scale gravity experiments, this approach has set independent limits and can aid the investigation of the geometric structure of the Universe.
@@
Figure 3: (a) Shows the changes in the behavior of the shadow radius with the variation of the $\mathrm{AdS_{5}}$ radius $\ell$. In turn, (b) we found bounds from the direct comparison between the results obtained by our theoretical analysis with the EHT experimental data.
IV Final Remarks
@@
Recently, higher-dimensional theories have drawn increasing attention. As has been pointed out, the seek for evidence of large extra dimensions has been performed in several experimental researches, making its detection a challenging problem in itself. By investigating the null geodesic curves, we find theoretical prediction for the black hole shadow radius, aiming to search for traces of extra dimensions. From recent measurements of the shadow image of Sgr A*—a supermassive black hole in the center of our Galaxy—we estimate deviations owing to an extra dimension in the RSII scenario on the theoretical shadow radius predicted. Given the inaccuracy involving the measured spin of Sgr A*, which may even be zero but still can provide compatible results with recent measures (Fragione:2020khu, ), we explore the possibility of describing them as a nonrotating RSII black hole. In this case, we have considered the Garriga-Tanaka solution that takes a classical approach to linearized braneworld gravity and describes a black hole solution in the weak field limit.
@@
As we saw, the curvature radius of the $\mathrm{AdS_{5}}$ spacetime is positively correlated to the shadow radius so that the increasing/decreasing in one leads to correspondingly increasing/decreasing in the other. Furthermore, we found that $\ell=0$, i.e., in the absence of new physics coming from extra dimension, one recovers the standard estimates expected for the Schwarzschild background. From recent observational data of the shadow of Sgr A* released by the EHT collaboration, one may impose constraints on the curvature radius of $\mathrm{AdS_{5}}$ spacetime. Therefore, by considering the Keck and VLTI estimates, we find more stringent limits on the $\mathrm{AdS_{5}}$ curvature radius, $\ell_{\left(\text{Keck}\right)}\lesssim 3.8\times 10^{-2}\,\mathrm{AU}$ and $\ell_{\left(\text{VLTI}\right)}\lesssim 1.6\times 10^{-2}\,\mathrm{AU}$ (at $68\%$ CL), than previously obtained for similar physical systems, to the best of our knowledge. Besides, a nonzero predicted curvature radius leads to agreement with the observed shadow radius at the $68\%$ confidence interval.
@@
We would like to highlight that, although our analysis provides weaker bounds than those obtained from collider experiments or sub-mm precision gravity tests, which have set limits on the non-Newtonian force, in this study, we have provided conservative bounds since we investigated the shadow radius in a nonrotating RSII-type black hole background. Further, the bounds we reported were obtained independently and in a manner complementary to the previously mentioned studies. Finally, the newly begun era of observational astronomy opens several avenues for future research that might be explored for testing fundamental physics—and so getting new constraints on the $\mathrm{AdS_{5}}$ curvature radius, for instance—or even searching eventual deviations on GR, shedding light into seeking physics beyond the SM.
@@ Acknowledgements.
We thank the referee for the valuable comments and suggestions that have enhanced this work. We would like to thank CNPq, CAPES, and CNPq/PRONEX/FAPESQ-PB (Grant No. $165/2018$), for partial financial support. F.A.B. acknowledges support from CNPq (Grant No. 309092/2022-1). A.S.L. acknowledges support from CAPES (Grant No. 88887.800922/2023-00). J.A.V.C. would like to thank FAPESQ-PB/CNPq (Grant No. 077/2022) for financial support.